# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.4 Proof of Theorem 3
$\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}\leq T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}+T^{2}(T-p+1)\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{F}.$
Result 11 (Auxiliary recurrences).
We have,
  1. 1.
    $\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{F}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}$
  2. 2.
    $\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{F}+\|\bm{W}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{W}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{F}$
  3. 3.
    $\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{F}\leq\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}}$
  4. 4.
    $\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{F}(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
    $\quad+2\|\bm{C}_{1,j}^{-1}\bm{U}_{1,j}\bm{V}_{1,j}-\bm{C}_{2,j}^{-1}\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{F}$
Lemma 12 (Bounding first term of Lemma 10).
The bounds in terms of $\bm{C}^{-1}_{k}=\bm{\Omega}+\sum_{j=1}^{k}\bm{L}_{j}\bm{L}_{j}^{T},\bm{U}_{k},\bm{V}_{k}$’s,
$\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{F}$
$\quad\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\sum_{j=0}^{p-1}\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F},$ (A.1)
@@
$\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}$
$\quad\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+p\sum_{k=1}^{p}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}M_{2,U})\big{]}$ (A.2)
where $M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p}$, $M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}},M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}$ for $\ell\in\{1,2\}$
Lemma 13 (Bounding second term of Lemma 10).
The final bounds for autoregressive coefficients,
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{F}$
$\quad\leq 2M_{1,U}M_{1,V}M_{1,C,U,V}\sum_{k=0}^{p}\|\bm{C}^{-1}_{1,k}-\bm{C}^{-1}_{2,k}\|^{2}_{F}+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}$
$\qquad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{F}M_{2,U}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}.$ (A.3)
where $M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}},M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}$, $M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p}$.
The following steps will transform the bounds for $\bm{U}_{j}$’s and $\bm{V}_{j}$’s to $\bm{L}_{j}$’s and $\bm{K}_{j}$’s.
@@
$\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{F}$
$\leq r\|\bm{K}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{K}_{j,1}^{\mathrm{T}}\bm{D}_{j-1,1}^{-1}\bm{K}_{1,1})^{-1}\|_{\mathrm{op}}\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-1,2}^{-1}\bm{K}_{j,2})^{-1}\|_{\mathrm{op}}\{\|\bm{K}_{j,1}\|_{\mathrm{op}}\|\bm{K}_{j,2}\|_{\mathrm{op}}\|\bm{D}_{j-1,1}^{-1}-\bm{D}_{j-1,2}^{-1}\|_{\mathrm{op}}$
$\quad+|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}}(\|\bm{D}_{j-1,1}^{-1}\|_{\mathrm{op}}\|\bm{K}_{j,1}\|_{\mathrm{op}}+\|\bm{D}_{j-1,2}^{-1}\|_{\mathrm{op}}\|\bm{K}_{j,2}\|_{\mathrm{op}})\}$
$\quad+r\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-1,2}^{-1}\bm{K}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}}$
$\|\bm{U}_{j,1}-\bm{U}_{j,2}\|^{2}_{\mathrm{op}}$
$\quad\leq\|\bm{C}_{j-1,1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1})^{-1}\|_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1}\|_{\mathrm{op}}\|\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1}$
$\qquad-\bm{L}^{\mathrm{T}}_{2}\bm{C}_{j-1,2}\bm{L}_{j,2}\|_{\mathrm{op}}+\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{C}_{j-1,2}\bm{L}_{j,2}\|^{2}_{\mathrm{op}}$
A.1 Proofs of the Auxiliary Lemmas and Resutls
@@
$\bm{M}={\begin{bmatrix}\bm{A}&\bm{C}^{\mathrm{T}}\\ \bm{C}&\bm{B}\end{bmatrix}}={\begin{bmatrix}\bm{I}_{p}&\bm{0}\\ \bm{C}\bm{A}^{-1}&\bm{I}_{q}\end{bmatrix}}{\begin{bmatrix}\bm{A}&\bm{0}\\ \bm{0}&\bm{B}-\bm{C}\bm{A}^{-1}C^{\mathrm{T}}\end{bmatrix}}{\begin{bmatrix}\bm{I}_{p}&\bm{A}^{-1}\bm{C}^{\mathrm{T}}\\ \bm{0}&\bm{I}_{q}\end{bmatrix}}.$
@@
The operator norm of $\bm{P}$ is 1 and $\bm{D}_{j}=\bm{\Gamma}(0)-\bm{\kappa}_{j}^{\mathrm{T}}\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}$.
@@
by applying the first inequality recursively as $\|\bm{D}_{j}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{p}^{-1}\|^{2}_{\mathrm{op}}$, $\bm{C}_{p}^{-1}=\bm{\Omega}+\sum_{k=1}^{p}\bm{L}_{K}\bm{L}_{k}^{\mathrm{T}}$.
Thus, $\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{F}\leq\|\bm{Q}_{1}-\bm{Q}_{2}\|_{\mathrm{F}}+(\|\bm{Q}_{1}\|^{2}_{\mathrm{op}}+\|\bm{Q}_{2}\|^{2}_{\mathrm{op}})\|\bm{P}_{1}-\bm{P}_{2}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{F}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}$.
Since, $\|\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\leq 1$, we have $\|\bm{\kappa}_{j}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{j-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{j-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Gamma}(0)\|^{2}_{\mathrm{op}}$
@@
There are $T$ many diagonal blocks of $\bm{\Upsilon}_{1}(0)$ and $\bm{\Upsilon}_{2}(0)$ in $\bm{\Upsilon}_{1,T}$ and $\bm{\Upsilon}_{2,T}$, respectively. Thus $\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\leq\frac{1}{T}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{F}$ and $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}$
$\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{op}}$. Thus, $\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}\leq\frac{\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}}{1-\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}}$ for small $\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}$.
Hence, $\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{F}\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}\frac{\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}}{1-\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}}$. This completes the proof.

### v2

@@ A.4 Proof of Theorem 3
The quantity $\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ is bounded by
$T^{2}\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}+T^{2}(T-p+1)\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}\|\bm{\kappa}_{1,p}\bm{\Upsilon}_{1,p-1}^{-1}-\bm{\kappa}_{2,p}\bm{\Upsilon}_{2,p-1}^{-1}\|^{2}_{\mathrm{F}}.$
Lemma 11 (Auxiliary recurrences).
The following bounds hold:
$\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{\mathrm{F}}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})$
×$\qquad\times\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}};$ (A.1)
$\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{\mathrm{F}}$
$\quad+\|\bm{W}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{W}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{\mathrm{F}};$ (A.2)
$\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}};$ (A.3)
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$
×$\qquad\times(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
$\quad+2\|\bm{C}_{1,j}^{-1}\bm{U}_{1,j}\bm{V}_{1,j}-\bm{C}_{2,j}^{-1}\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{\mathrm{F}}.$ (A.4)
Lemma 12 (Bounding the first term in Lemma 10).
The quantity $\|\bm{\Upsilon}_{1,p-1}-\bm{\Upsilon}_{2,p-1}\|^{2}_{\mathrm{F}}$ is bounded by
$\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}+\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ (A.5)
$\qquad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\sum_{j=0}^{p-1}\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}},$
and
$\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq{p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}}$
$\quad+p\sum_{k=1}^{p}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{F}}$
$\qquad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})$
×$\qquad\qquad\times(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}M_{2,U})\big{]},$ (A.6)
where, for $\ell\in\{1,2\}$,
$M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p},\quad M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}},M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}.$
Lemma 13 (Bounding the second term in Lemma 10).
@@
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$ $\leq{2M_{1,U}M_{1,V}M_{1,C,U,V}\sum_{k=0}^{p}\|\bm{C}^{-1}_{1,k}-\bm{C}^{-1}_{2,k}\|^{2}_{\mathrm{F}}}$
$\quad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}$
$\quad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{\mathrm{F}}M_{2,U}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}},$ (A.7)
@@
$M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}},\quad M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}},\quad M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p}.$
The following steps transform the bounds for $\bm{U}_{j}$’s and $\bm{V}_{j}$’s to $\bm{L}_{j}$’s and $\bm{K}_{j}$’s.
@@
The quantity $\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{\mathrm{F}}$ is bounded by
$r\|\bm{K}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{K}_{j,1}^{\mathrm{T}}\bm{D}_{j-1,1}^{-1}\bm{K}_{1,1})^{-1}\|_{\mathrm{op}}\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-1,2}^{-1}\bm{K}_{j,2})^{-1}\|_{\mathrm{op}}$
×$\quad\times\{\|\bm{K}_{j,1}\|_{\mathrm{op}}\|\bm{K}_{j,2}\|_{\mathrm{op}}\|\bm{D}_{j-1,1}^{-1}-\bm{D}_{j-1,2}^{-1}\|_{\mathrm{op}}$
$+|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}}(\|\bm{D}_{j-1,1}^{-1}\|_{\mathrm{op}}\|\bm{K}_{j,1}\|_{\mathrm{op}}+\|\bm{D}_{j-1,2}^{-1}\|_{\mathrm{op}}\|\bm{K}_{j,2}\|_{\mathrm{op}})\}$
$+r\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-1,2}^{-1}\bm{K}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}},$
and $\|\bm{U}_{j,1}-\bm{U}_{j,2}\|^{2}_{\mathrm{op}}$ is bounded by
$\|\bm{C}_{j-1,1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1})^{-1}\|_{\mathrm{op}}$
×$\quad\times\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1}\|_{\mathrm{op}}\|\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{L}^{\mathrm{T}}_{2}\bm{C}_{j-1,2}\bm{L}_{j,2}\|_{\mathrm{op}}$
$+\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{C}_{j-1,2}\bm{L}_{j,2}\|^{2}_{\mathrm{op}}.$
A.1 Proofs of the Auxiliary Lemmas
@@
$\bm{M}={\begin{bmatrix}\bm{A}&\bm{C}^{\mathrm{T}}\\ \bm{C}&\bm{B}\end{bmatrix}}={\begin{bmatrix}\bm{I}_{p}&\bm{0}\\ \bm{C}\bm{A}^{-1}&\bm{I}_{q}\end{bmatrix}}{\begin{bmatrix}\bm{A}&\bm{0}\\ \bm{0}&\bm{B}-\bm{C}\bm{A}^{-1}C^{\mathrm{T}}\end{bmatrix}}{\begin{bmatrix}\bm{I}_{p}&\bm{A}^{-1}\bm{C}^{\mathrm{T}}\\ \bm{0}&\bm{I}_{q}\end{bmatrix}},$
we have
@@
The operator norm of $\bm{P}$ is 1 and $\bm{D}_{j}=\bm{\Gamma}(0)-\bm{\kappa}_{j}^{\mathrm{T}}\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}$, giving
@@
by applying the first inequality recursively since $\|\bm{D}_{j}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{p}^{-1}\|^{2}_{\mathrm{op}}$ and $\bm{C}_{p}^{-1}=\bm{\Omega}+\sum_{k=1}^{p}\bm{L}_{K}\bm{L}_{k}^{\mathrm{T}}$. Thus,
$\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{Q}_{1}-\bm{Q}_{2}\|_{\mathrm{F}}+(\|\bm{Q}_{1}\|^{2}_{\mathrm{op}}+\|\bm{Q}_{2}\|^{2}_{\mathrm{op}})\|\bm{P}_{1}-\bm{P}_{2}\|^{2}_{\mathrm{F}}$
$\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{\mathrm{F}}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}.$
Thus we have
$\|\bm{\kappa}_{j}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{j-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{j-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Gamma}(0)\|^{2}_{\mathrm{op}},$
since $\|\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\leq 1$, ∎
@@
There are $T$ many diagonal blocks of $\bm{\Upsilon}_{1}(0)$ and $\bm{\Upsilon}_{2}(0)$ in $\bm{\Upsilon}_{1,T}$ and $\bm{\Upsilon}_{2,T}$, respectively. Thus, $\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}\leq T^{-1}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{F}}$ and
$\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}$ $\leq\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1,T}-\bm{\Upsilon}_{2,T}\|^{2}_{\mathrm{op}},$
which leads to the bound
$\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}\leq\frac{\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}}{1-\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}}$
whenever $\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}$ is sufficiently small. Hence,
$\|\bm{\Omega}_{1}-\bm{\Omega}_{2}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{2}(0)^{-1}\|^{2}_{\mathrm{op}}$
$\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}\frac{\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}}{1-\|\bm{\Upsilon}_{1}(0)^{-1}\|^{2}_{\mathrm{op}}\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}}}.$
This completes the proof. ∎